# Systematic Uncertainties in the Measurement of Neutron lifetime Using Lunar Prospector Neutron Spectrometer

### v2

@@ IV Computing Neutron Lifetime
We model the neutron propagation to compute the HeCd and HeSn detector count rates, and calculate thermal neutrons. A simple reduced $\chi^{2}$ statistic is used to find the best-fit neutron lifetime using thermal neutrons as the model parameter. We use $\approx$50 different neutron lifetime values in a broad prior ranging $\tau_{n}=(540,1680)$ s. This analysis is repeated for each cases of (a) 20 resolution, 5 resolution, 5 compositions reported in Wilson et al. (2021) to study the effect of surface composition; and (b) 20 resolution at constant equatorial temperature, and latitude dependent sub-surface temperature model to study the effect of surface temperature on the measurement of neutron lifetime.
@@
It can be noted from Figure 4 that the model is offset from the data, predominantly in the lower altitude regimes. To capture the shortcomings of the model, we scale the thermal neutrons by a constant factor to minimize the discrepancy. The scale factor is constant for a given comparison between model and data, but is recalculated for every $\tau_{n}$ value. This approach of accounting for the model misfit has also been used by Wilson et al. (2021), however the value of the chosen multiplicative factor has not been reported. Figure 8 shows the scaling factor computed for each of the cases capturing lunar composition and temperature effects. The five compositions used by Wilson et al. (2021) need the lowest scaling, whereas a model with realistic conditions that capture the complex lunar compositions and temperature effects needs a slightly higher scaling factor. This indicates that a more realistic lunar model of neutron propagation fails to capture the neutron spectra that closely agrees with the LP data. We also note that, using analysis descriptions and assumptions given in Wilson et al. (2021), we fail to retrieve the reported neutron lifetime of $\tau_{n}=887\pm 14_{stat}$ s. The notable difference between the two analyses is (i) we use MCNP 6.2, where as MCNPX has been used in Wilson et al. (2021). (ii) We do not apply temporal offset to the spacecraft altitude and velocity to account for the mismatch in the ephemerides and GRS data. The analysis in Wilson et al. (2021) found that the time of the ephemeris did not agree with the time of the count data and required a 900 s offset. We do not find evidence of this offset and suspect that we are using a different version of the ephemerides data. (iii) Wilson et al. (2021) analysis is confined to those observations where LP’s longitude was greater than 180$\circ$ which removed the eastern lunar maria from the study as the residuals in this region were found to be anomalously large. This suggests that the complex composition of this region has not been fully incorporated into their model. This work in contrast uses the full orbital insertion data that captures the complex lunar neutron emission fields.
Table 3: Computed $\tau_{n}$ values for each of the cases capturing lunar surface composition and temperature effects.
@@
Five compositions Latitude independent 739.64 $\pm$ 10.83
@@
20 Latitude independent 738.63 $\pm$ 10.75
20 Latitude dependent 767.33 $\pm$ 11.17
5 Latitude dependent 777.59 $\pm$ 11.71
@@
Figure 8: Left: Correction factor to account for the model and data offset. Each scaling factor is used as a single multiplicative factor to the thermal neutrons, which is used in reduced $\chi^{2}$ analysis. The five composition used by Wilson et al. (2021) needs the lowest scaling, where as a realistic model capturing complex lunar terrains and latitude dependent temperature effects needs slightly bigger scaling factor. Right: Reduced $\chi^{2}$ analysis for the each of the test cases and their corresponding minimum value shown in dashed lines.
@@ V Future Work
This work along with recent space-based measurements demonstrate the efficacy of the technique with limited non-optimized data. As these measurements are limited by systematic errors, a competitive measurement using space-based technique demands more data using neutron detectors and orbits specific for neutron lifetime measurements. Currently there are no planned orbital lunar neutron measurements on missions in development or opportunities to propose such a mission. Small missions like LunaH-Map, a cubesat launched in November 2022 on the Space Launch System Artemis-1 mission, could provide an opportunity to demonstrate the collection of neutron data from elliptical lunar orbits and the ability of space-based measurements to quantify the neutron lifetime Hardgrove et al. (2020). Although LunaH-Map was not designed for neutron lifetime measurements, the spacecraft’s neutron spectrometer collected lunar neutron and gamma-ray data demonstrating the low-cost pathfinder mission concept for lunar exploration and science Hardgrove et al. (2023). Additional work on space-based neutron lifetime measurements will focus on the selection of an appropriate planetary body and orbital parameters for spacecraft insertion among Earth, Moon and Venus, as well as the appropriate instrumentation (i.e. selection of neutron detector). A dedicated mission for measuring the neutron lifetime would have specific orbital requirements for the spacecraft to obtain a larger data-set and achieve a higher confidence measurement of neutron lifetime.
The preliminary predicted statistical precision, mission and orbit requirements are laid out in Lawrence et al. (2021), where Venus is determined to be the best planetary body due to higher statistical precision and simpler surface composition. However, the study lacks study of uncertainties in the neutron lifetime due to its atmospheric composition and temperatures. While Earth is also determined to be a good choice due to ease of orbit insertion and relatively inexpensive mission costs, lunar orbital and landed measurements are equally viable. Although this work quantifies the effects of lunar surface composition and temperature, a detailed study of systematic uncertainties for Venus, Earth and Moon are necessary for a dedicated neutron lifetime mission.
@@ VI Conclusion
[…]

### v3

@@ IV Computing Neutron Lifetime
We model the neutron propagation to compute the HeCd and HeSn detector count rates, and calculate thermal neutrons. A simple $\chi^{2}$ statistic is used to find the best-fit neutron lifetime using thermal neutrons as the model parameter. We use $\approx$50 different neutron lifetime values in a broad prior ranging $\tau_{n}=(540,1680)$ s. This analysis is repeated for each cases of (a) 20 resolution, 5 resolution, 5 compositions reported in Wilson et al. (2021) to study the effect of surface composition; and (b) 20 resolution at constant equatorial temperature, and latitude dependent sub-surface temperature model to study the effect of surface temperature on the measurement of neutron lifetime.
@@
It can be noted from Figure 4 that the model is offset from the data, predominantly in the lower altitude regimes. To capture the shortcomings of the model, we scale the thermal neutrons by a constant factor to minimize the discrepancy. The scale factor is constant for a given comparison between model and data, but is recalculated for every $\tau_{n}$ value. This approach of accounting for the model misfit has also been used by Wilson et al. (2021), however the value of the chosen multiplicative factor has not been reported. Figure 8 shows the scaling factor computed for each of the cases capturing lunar composition and temperature effects. The five compositions used by Wilson et al. (2021) need the lowest scaling, whereas a model with realistic conditions that capture the complex lunar compositions and temperature effects needs a slightly higher scaling factor. This indicates that a more realistic lunar model of neutron propagation fails to capture the neutron spectra that closely agrees with the LP data. We also note that, with our best understanding of the analysis descriptions and assumptions given in Wilson et al. (2021) along with details of the detector response shared through private communications, we fail to retrieve the reported neutron lifetime of $\tau_{n}=887\pm 14_{stat}$ s. The notable difference between the two analyses is (i) we use MCNP 6.2, where as MCNPX has been used in Wilson et al. (2021). (ii) We do not apply temporal offset to the spacecraft altitude and velocity to account for the mismatch in the ephemerides and GRS data. The analysis in Wilson et al. (2021) found that the time of the ephemeris did not agree with the time of the count data and required a 900 s offset. We do not find evidence of this offset and suspect that we are using a different version of the ephemerides data. (iii) Wilson et al. (2021) analysis is confined to those observations where LP’s longitude was greater than 180 which removed the eastern lunar maria from the study as the residuals in this region were found to be anomalously large. This suggests that the complex composition of this region has not been fully incorporated into their model. These choices lead to an overall lower scaling factor used to offset the model, resulting in better agreement with LPNS data. This work in contrast uses the full orbital insertion data that captures the complex lunar neutron emission fields, and shows the effects of choice of compositional and temperature models on the resulting neutron lifetime measurement.
[…]
@@
Five compositions Latitude independent 739.6 $\pm$ 10.8
@@
20 Latitude independent 738.6 $\pm$ 10.8
20 Latitude dependent 767.3 $\pm$ 11.2
5 Latitude dependent 777.6 $\pm$ 11.7
@@
By explicitly quantifying the impact of subsurface temperature models and compositional maps on the derived neutron lifetime, we demonstrate that these two systematics alone introduce significant variance, considerably larger than the estimated systematic uncertainties derived in Wilson et al. (2021), yet still fall well short of explaining the full discrepancy between our results and those previously reported. While Wilson et al. (2021) applied an empirical scaling factor to reconcile their model with observational data, we find that this approach is insufficient when applied to our independent analysis, further suggesting that additional unaccounted-for systematics are influencing these measurements. These findings highlight a broader issue: that current space-based measurements of the neutron lifetime, particularly those relying on legacy datasets such as from Lunar Prospector, remain subject to large and poorly constrained uncertainties.
We show the results from the four models alongside the reported neutron lifetime for bottle, beam and recent space-based measurements using MESSENGER flybys of Venus and Mercury; and LP data in Figure 9. Space-based measurements are farther from the current laboratory measurements, however these are intended to serve as a demonstration of the technique using limited non-optimized data. Although surface composition and latitudinal temperature are expected to be significant sources of systematic uncertainties in the measurement, this work is not a comprehensive study of all the systematics. Additional systematics that may not be fully removed by the use of a scaling factor include modeling of the detector response that requires accurate calibration, contributions from uncertainties of simulation physics models, and uncertainties in the GCR spectrum. This work serves to illustrate the effects of choice of the compositional and temperature models on the computed neutron lifetime. A dedicated and rigorous systematic uncertainty analysis, including precise detector calibration and independent lunar surface modeling is essential for future competitive neutron lifetime measurements from space-based neutron data. Such analysis warrants a separate, focused investigation that is beyond the scope of this work.
Figure 8: Left: Correction factor to account for the model and data offset. Each scaling factor is used as a single multiplicative factor to the thermal neutrons, which is used in $\chi^{2}$ analysis. The five composition used by Wilson et al. (2021) needs the lowest scaling, where as a realistic model capturing complex lunar terrains and latitude dependent temperature effects needs slightly bigger scaling factor. Right: $\chi^{2}$ analysis for the each of the test cases and their corresponding minimum value shown in dashed lines.
@@ V Future Work
This work along with recent space-based measurements demonstrate the efficacy of the technique with limited non-optimized data. As these measurements are limited by systematic errors, a competitive measurement using space-based technique demands more data using neutron detectors and orbits specific for neutron lifetime measurements. Currently there are no planned orbital lunar neutron measurements on missions in development or opportunities to propose such a mission. Small missions like LunaH-Map, a cubesat launched in November 2022 on the Space Launch System Artemis-1 mission, could provide an opportunity to demonstrate the collection of neutron data from elliptical lunar orbits and the ability of space-based measurements to quantify the neutron lifetime Hardgrove et al. (2020). Although LunaH-Map was not designed for neutron lifetime measurements, the spacecraft’s neutron spectrometer collected lunar neutron and gamma-ray data demonstrating the low-cost pathfinder mission concept for lunar exploration and science Hardgrove et al. (2023). Additional work on space-based neutron lifetime measurements will focus on the selection of an appropriate planetary body and orbital parameters for spacecraft insertion among Earth, Moon and Venus, as well as the appropriate instrumentation (i.e. selection of neutron detector). A dedicated mission for measuring the neutron lifetime would have specific orbital requirements for the spacecraft to obtain a larger data-set and achieve a higher confidence measurement of neutron lifetime. The overall increase in data volume, better detector response simulations and modeling of the neutron count rates will help understand and eliminate large systematics, potentially leading to a competitive space-based neutron lifetime measurement.
The preliminary predicted statistical precision, mission and orbit requirements are laid out in Lawrence et al. (2021), where Venus is determined to be the best planetary body due to higher statistical precision and simpler surface composition. However, the work lacks study of uncertainties in the neutron lifetime due to its atmospheric composition and temperatures. While Earth is also determined to be a good choice due to ease of orbit insertion and relatively inexpensive mission costs, lunar orbital and landed measurements are equally viable. A detailed study of systematic uncertainties for Venus, Earth and Moon are necessary for a dedicated neutron lifetime mission. This work quantifies the effects of lunar surface composition and temperature but identifies significant unaccounted-for systematics that may explain the large discrepancy from the PDG value. In light of detector calibration limitations, spacecraft geometry effects (e.g., unmodeled shadowing), and the suboptimal nature of the dataset used, future efforts may benefit from the use of improved instrumentation, tailored orbital geometries, more homogeneous targets, and efforts to further evaluate systematic uncertainties that may offer a more robust pathway toward a competitive neutron lifetime measurement with space-based observations.
@@ VI Conclusion
We used the HeCd and HeSn detector data during Lunar Prospector’s orbit insertion period, where it transitioned from a highly elliptical orbit to a circular one where it was parked for its mission, and computed the neutron lifetime. We used MCNP to get the lunar neutron current, and applied particle kinematics to model the neutron propagation before it hits the LP detector. Thermal neutrons ($E<0.4$ eV) were used to calculate the neutron lifetime using a $\chi^{2}$ fit, after a scalar multiplicative correction to account for the misfit between data and model. We also show the effects of precise lunar surface composition and temperature. We find $\tau_{n}$ to be dependent on surface composition and temperatures, varying from $739.6\pm 10.8$ s, for a simple five compositions re-binned to 2 resolution and assuming a constant equatorial lunar surface temperature, $738.6\pm 10.8$ s for 20 resolution map and assuming constant temperature, $767.3\pm 11.2$ s assuming latitude dependent temperature, and $777.6\pm 11.7$ s for 5 resolution map and assuming latitude dependent temperature. Our analysis underscores the challenges in replicating the neutron lifetime results reported by Wilson et al. (2021), even after extensive efforts to align methodological assumptions and data treatments. The reported measurements here are not yet competitive with the current PDG value pointing to additional systematics that are not yet accounted for. However, this work serves to illustrate the effect on the neutron lifetime due to choice of temperature and compositional models. We estimate the effect on neutron lifetime due to choice of temperature model to be 28.7 $\pm$ 15.5 s, and choice of compositional map (for 20 and 5 maps) to be 10.3 $\pm$ 12.2 s. This analysis was done with a limited non-optimized data, thus showing the potential of such space-based technique for future precise measurements of neutron lifetime.